The lowest limit of MgSO4 that can be detected by taste in drinking water is 0.400g/L. What is the molar concentration of MgSO4?
1
Expert's answer
2019-03-21T06:25:45-0400

Mr (MgSO4) = 120.37 g/mol

n = m/Mr = 0.4/120.37 = 0.003 mol

Molar concentration - 0.003 M